# Pi in the Sky: Neutron Stars with Exceptionally Light QCD Axions

### v1

@@ 3.2 Axion condensation in Chiral Effective Field Theory
ChiEFT based calculations reported in Refs. Kaiser:2007nv ; Plohl:2007dp ; Kruger:2013iza find that $R(n_{B})/R_{0}(n_{B})>1$ over the entire range of densities where the ChiEFT is expected to be useful. Thus, the interaction energy increases with decreasing pion mass in the vicinity of $m_{\pi}=m^{\rm phys}_{\pi}$ in SNM and PNM. However, the study in Ref. Kaiser:2007nv , which was restricted to the case of SNM, found that the ratio $R(n_{B})/R_{0}(n_{B})<1$ for smaller $m_{\pi}$. Their calculations indicate that $R_{0}(n_{\rm sat})-R(n_{\rm sat})\simeq 0.1$ at $m_{\pi}=70$ MeV and $R_{0}(n_{\rm sat})-R(n_{\rm sat})\simeq 0.3$ in the chiral limit. Together, these results imply that the interaction energy increases with decreasing $m_{\pi}$ for $m_{\pi}\simeq m^{\rm phys}_{\pi}$ and then decreases rapidly for smaller values of $m_{\pi}$. The microscopic calculations of the energy per particle of PNM and SNM as a function of $m_{\pi}$ that we shall describe in the following show a similar trend.
@@
Figure 2: Difference in the energy per particle (17) for the N2LO sim potentials Carlsson:2015vda with two momentum cutoffs $\Lambda$ at $n_{B}=n_{\rm sat}$ as a function of $m_{\pi}$: $\Lambda=500$ MeV (left panel) and $\Lambda=450$ MeV (right panel), as annotated. In these potentials, we modify $m_{\pi}$, $f_{\pi}$, and the LO LEC $\tilde{C}_{{}^{1}S_{0}}$ (different line styles), whereas $g_{A}$ is kept at its physical value. The bands depict the variation of the LO LEC $\tilde{C}_{{}^{1}S_{0}}$ according to Eq. (19). The results were obtained at the Hartree-Fock level (green bands) and using second-order (blue bands) and third-order MBPT (red bands). See the main text for details.
Figure 2 shows the difference in the PNM energies per particle
$\Delta E(n_{\rm sat},m_{\pi})=E(n_{\rm sat},m_{\pi})-E(n_{\rm sat},m^{\rm phys}_{\pi})$ (17)
@@
for two chiral potentials at $n_{B}=n_{\rm sat}$ as a function of $m_{\pi}$ (see annotations). Here, we use the family of bare N2LO potentials with chiral NN and 3N forces developed in Ref. Carlsson:2015vda . These nonlocal potentials are called the sim potentials because they were simultaneously fit to scattering and bound-state observables in the pion-nucleon, NN, and few-nucleon sectors. Specifically, we use the two N2LO potentials with the NN and 3N momentum cutoffs $\Lambda=500$ (left panel) and $450$ MeV (right panel); both constrained by NN scattering data at laboratory energies $T_{\mathrm{lab}}\leqslant 290$ MeV Carlsson:2015vda . To obtain the PNM energy per particle, we apply the Monte Carlo framework for MBPT nuclear matter calculations developed in Ref. Drischler:2017wtt . The green bands in Fig. 2 depict the Hartree-Fock reference energy and the blue and red bands the MBPT results at second and third order, respectively. All MBPT calculations are performed at the normal-ordered two-body level. More details on MBPT for nuclear matter can be found in Ref. Drischler:2021kxf . This MBPT framework uses the operatorial definition of the chiral potentials, allowing us to straightforwardly vary the values of physical constants, such as the pion mass, and low-energy couplings, and more. Not all chiral potentials are given in their operatorial form; e.g., when evolved to lower resolution scales using renormalization group (RG) methods, their operatorial definition is usually unknown. For the solid lines in Fig. 2, we keep the values of the low-energy couplings at their best-fit values determined in Ref. Carlsson:2015vda and vary only the pion mass, nucleon mass, and the pion decay constant. The $m_{\pi}$ dependence of the nucleon mass was discussed in Sec. 2.1, and the $m_{\pi}$ dependence of the pion decay constant is given by
@@
with its value in the chiral limit $f=86.2$ MeV and the LEC $\bar{l}_{4}=4.3$ Berengut:2013nh . The pion-mass dependence of the axial coupling constant $g_{A}$ is relatively weak in the regime considered here, as can be seen in Figure 1 in Ref. Berengut:2013nh , and thus we keep $g_{A}=1.29$ at the physical value. We use these potentials to identify overall trends and emphasize the need to explore the pion-mass dependence of a wider range of chiral potentials in the future.
@@
where $f(\theta)=\left(m_{\pi}(\theta)/m^{\rm phys}_{\pi}\right)^{2}$, $\eta=D_{2}(m^{\rm phys}_{\pi})^{2}/\tilde{C}_{{}^{1}S_{0}}^{\rm fit}$ and $\tilde{C}_{{}^{1}S_{0}}^{\rm fit}$ is the best-fit value in each isospin channel obtained in Ref. Carlsson:2015vda . Here, $D_{2}$ is the LEC discussed in Ref. Kaplan:1998tg ; Beane:2002vs that captures the pion-mass dependence at NLO. We choose the two values $\eta=\pm 0.1$ to set the limits of this variation and adopt $m^{\rm phys}_{\pi}=138.04$ MeV for the physical pion mass. The dotted and dash-dotted lines in Fig. 2 depict the results for the PNM energy per particle with the modified LO coupling constant (19); the solid lines correspond to $\eta=0$ and thus $\tilde{C}_{{}^{1}S_{0}}(\theta)\equiv\tilde{C}_{{}^{1}S_{0}}$. Increasing the strength of $\tilde{C}_{{}^{1}S_{0}}(\theta)$, corresponding to $\eta=-0.1$, reduces $\Delta E(n_{\rm sat},m_{\pi})$, whereas decreasing the strength, corresponding to $\eta=+0.1$, increases $\Delta E(n_{\rm sat},m_{\pi})$.
Several important insights can be drawn from the results in Fig. 2. MBPT appears to converge more slowly with decreasing pion mass, which could be further investigated using a Weinberg eigenvalue analysis Hoppe:2017lok . Although the leading Hartree-Fock energies indicate a systematic increase in interaction energy with $m_{\pi}$ across the range shown in Fig. 2, the second-order correction reverses this trend. The third-order correction is modest and of the same sign as the second-order correction. For relatively small reductions of the pion mass ($m_{\pi}\gtrsim 120$ MeV for $\Lambda=500$ MeV and $m_{\pi}\gtrsim 110$ MeV for $\Lambda=450$ MeV), the energy difference (17) is positive. The sensitivity of the results at second and third order to the cutoff variation is larger than that of the LEC variation. These general trends are not altered by varying the momentum cutoff over the range shown in Fig. 2. However, the rather large dependence on the momentum cutoff is disconcerting and warrants a careful inclusion of the $m_{\pi}$ dependence of the short-range components of the force. As discussed earlier, we have studied the effect of the $m_{\pi}$ dependence of the leading LEC associated with scattering in the ${}^{1}S_{0}$ channel shows that the qualitative behavior is not changed. Based on these trends, we conclude that the interaction energy at $\theta=\pi$, where $m_{\pi}\approx 80$ MeV can be smaller than at $\theta=0$, where $m_{\pi}\approx 138$ MeV. However, the energy difference $\Delta E$ remains quite sensitive to the regularization scheme.
@@
3.3 Axion condensation in Relativistic Mean Field Theory
To derive concrete results for neutron stars, we use a model within RMFT (see Ref. Dutra:2014rmfreview for a review). The Lagrangian of this theory is given by the following, omitting terms that will be discarded after taking the meson fields to have their mean field values.
@@ 3.3 Axion condensation in Relativistic Mean Field Theory
$\mathcal{L}_{\sigma\omega\rho}$ contains all meson-meson couplings that are fitted along with the nucleon-meson couplings and meson masses to known properties of nuclear matter and $\tau_{3}$ is the third Pauli matrix in isospin space. We use the IUFSU parameter set from Ref. Agrawal:2012rmfmodel . The mean field values of the mesons and Fermi momenta of the nucleons and leptons are calculated by solving the Euler-Lagrange equations for each field and enforcing charge neutrality and $\beta$ equilibrium.
@@
For Scenario A, the simple case where only the nucleon masses are modified by finite $\theta$, Fig. 3 shows the energy per baryon, including axion energy and the reduction of the mass of the nucleons, of SNM, PNM, and beta equilibrated matter in this model. At fixed baryon density, the energy per baryon is minimized, leading to axion condensation due to the reduced nucleon mass. For small $\varepsilon$, the energy cost of having non-zero $\theta$ is small and axion condensation is favored at lower densities. Since the axions themselves contribute to the energy density, the energy per baryon of axion condensed matter will always diverge in the limit of zero density and a phase without axions will be favored. This has the effect of producing a minimum in the energy per baryon of axion condensed matter when $\varepsilon$ is sufficiently small and allows a stable zero pressure phase of nuclear matter at lower energy than normal nuclei. We will return to discuss these objects in Sec. 5.
@@
Of particular relevance to observations of neutron stars is the pressure at which neutron drip occurs. Neutron drip refers to the density at which it becomes energetically favorable for neutrons to come out of nuclei and form a background of dripped neutrons. As such, the neutron drip phase transition will occur when the baryon chemical potential is equal to the mass of the neutron (possibly modified by finite $\theta$) in the case where PNM is not self-bound. If the axion-condensed phase persists to the surface of the neutron star, but negative axion pressure is larger than the pressure of electrons at neutron drip density (nuclei make a small contribution to the pressure), the neutron drip phase transition will not be found in the neutron star and the surface of the star will have a thin domain wall covering either a phase of nuclei in a dripped neutron background or bulk nuclear matter.
Not included thus far is the pion mass dependence of the nuclear force itself. The $\theta$-dependence of the masses of the $\sigma$ and $\rho$ mesons have been calculated in Refs. HanhartPelaezRios2008 ; Pelaez:2010sigmamass . They find that, comparing $\theta=0$ and $\theta=\pi$, the $\sigma$ mass decreases by about 6% and the $\rho$ mass decreases by about 2%. At low density, the RMFT EOS is dominated by the $\sigma$ meson and its modification is much larger than that of the vector mesons so we focus on the effect of modifying the scalar interaction. As a simple ansatz for modifications to the nuclear force, we modify the $\sigma$ mass with a tunable parameter $d_{\sigma}$ as follows:
@@
where $g_{\sigma}$ is a parameter of the RMFT model and $\chi$ is the volume fraction of nuclei when in a heterogeneous phase and one in a homogeneous phase. At saturation density, $g_{\sigma}^{2}n_{\rm sat}/\sigma_{N}m_{\sigma}^{2}\simeq 11$ for the RMFT model we use so for Scenario B, axion condensation should be expected below $n_{\rm sat}$ for $\varepsilon\lesssim 0.8$ and in Scenario C, axion condensation should be expected below $n_{\rm sat}$ even for $\varepsilon=1$! The predictive power of this model is limited by the fact that we only consider isoscalar effects on the nuclear force. Axion condensation may also have an important impact on the symmetry energy of nuclear matter beyond the effect of $\Delta\sigma$. We expect this to have a more pronounced effect at high density where our model of the nucleon-axion interaction breaks down.
Figure 4 shows the kinetic and interaction energy per baryon for Scenarios A, B, and C at $\theta=\pi$. Since this only includes the energy from kinetic energy of baryons and nuclear forces, this quantity is independent of $\varepsilon$. Note that this differs from Fig. 3 in that the reduction of the nucleon masses is not included in this value.
@@
When $\varepsilon$ is small enough that axion condensed matter is favored down to zero pressure, there can be striking effects on the crust of the star. We construct an axion-condensed inner crust of nuclei in a background of dripped neutrons, including Coulomb and surface energies. Since the constraints we derive on axion parameter space are tied to the properties of the crust, we choose a surface tension $\sigma=0.75\,\mbox{MeV}/\mbox{fm}^{2}$ tuned to recreate neutron drip pressure at $\theta=0$ calculated with more detailed microphysics (see, e.g., Ref. BPS1971 ). To be self-consistent between the two phases, the phase transition will be calculated between the RMFT model in all crust phases for all values of $\theta$ rather than using a tabulated crust EOS. Where relevant, we compare quantitative results to the calculation of crust properties calculation in the BPS and NV EOSs BPS1971 ; NegeleVautherin1973 .
@@
Figure 5 shows a phase diagram for the RMFT model as a function of pressure and $\varepsilon$. In the blue-shaded region, $\theta=\pi$ is favored while in the unshaded region $\theta=0$ is favored. The cross-hatched regions contain the outer crust of nuclei and electrons and the diagonal-hatched regions contain the inner crust of nuclei, dripped neutrons, and electrons. Regions without hatching are composed of homogeneous nuclear matter. The diagonal colored lines show the domain wall pressure for an axion of a given mass with the specified $\varepsilon$ if the radius of the domain wall is 12 km. When the boundary region is treated as infinitely thin, this is the pressure difference across the domain wall. For small $\varepsilon$ where the axion condensed phase is favored for all densities, this acts as a cutoff pressure at the surface of the star; lower pressures can only exist in the finite size of the domain wall itself. Effects of the domain wall will be discussed in more detail in Sec. 4.2.
@@
Figure 5: Phase diagram for RMFT model with IUFSU parameter set and our three choices of $K_{\sigma}$. The colored diagonal lines show the domain wall pressure for $R=12$ km. Blue shaded: $\theta=\pi$. Unshaded: $\theta=0$. Cross-hatched: outer crust. Diagonal hatched: inner crust. No hatch: core.
@@
As $K_{\sigma}$ is increased from 00, the pressure of the PNM surrounding the nuclei in the inner crust decreases significantly and, as a result, the chemical potential corresponding to the crust-core transition (the boundary between the diagonal hatched and unhatched regions) decreases. At the same time, a more attractive nuclear interaction causes nuclei to become more neutron rich, increasing the electron chemical potential. An increased electron density corresponds to a larger pressure at neutron drip (the boundary between the diagonal hatched and cross hatched regions). For sufficiently large $K_{\sigma}$, if $\varepsilon$ is small enough for an axion condensed crust to be present, there will be a phase transition directly from bulk nuclear matter to a phase with nuclei and free electrons, with no dripped neutrons. This occurs because bulk matter is favored when the chemical potential is less than $m_{n}(\theta=\pi)+\Gamma_{n}(\theta=\pi)$ where $\Gamma_{n}(\theta)$ is the binding of PNM in the case where neutron matter is self-bound (Scenarios B and C) and 00 in the case where it is not.
3.4 EOS of axion condensed matter

### v2

@@ III.2 Axion condensation in Chiral Effective Field Theory
ChiEFT-based calculations reported in Refs. [69, 70, 71] find that $R(n_{B})/R_{0}(n_{B})>1$ over the entire range of densities where the ChiEFT is expected to be useful. Thus, the interaction energy increases with decreasing pion mass in the vicinity of $m_{\pi}=m^{\rm phys}_{\pi}$ in SNM and PNM. However, the study in Ref. [69], which was restricted to the case of SNM, found that the ratio $R(n_{B})/R_{0}(n_{B})<1$ for smaller $m_{\pi}$. Their calculations indicate that $R_{0}(n_{\rm sat})-R(n_{\rm sat})\simeq 0.1$ at $m_{\pi}=70$ MeV and $R_{0}(n_{\rm sat})-R(n_{\rm sat})\simeq 0.3$ in the chiral limit. Together, these results imply that the interaction energy increases with decreasing $m_{\pi}$ for $m_{\pi}\simeq m^{\rm phys}_{\pi}$ and then decreases rapidly for smaller values of $m_{\pi}$. The microscopic calculations of the energy per particle of PNM as a function of $m_{\pi}$ that we shall describe in the following show a similar trend.
@@
Figure 2: Difference in the energy per particle (17) for the N2LO sim potentials [72] with two momentum cutoffs $\Lambda$ at $n_{B}=n_{\rm sat}$ as a function of $m_{\pi}$: $\Lambda=500$ MeV (left panel) and $\Lambda=450$ MeV (right panel), as annotated. In these potentials, we modify $m_{\pi}$, $f_{\pi}$, and the LO LEC $\tilde{C}_{{}^{1}S_{0}}$ (different line styles), whereas $g_{A}$ is kept at its physical value. The bands depict the variation of the LO LEC $\tilde{C}_{{}^{1}S_{0}}$ according to Eq. (19). The results were obtained at the Hartree-Fock level (green bands) and using second-order (blue bands) and third-order MBPT (red bands). See the main text for details.
Figure 2 shows the difference in the PNM energies per particle,
$\Delta E(n_{\rm sat},m_{\pi})=E(n_{\rm sat},m_{\pi})-E(n_{\rm sat},m^{\rm phys}_{\pi}),$ (17)
@@
for two chiral potentials at $n_{B}=n_{\rm sat}$ as a function of $m_{\pi}$ (see annotations). Here, we use the family of bare N2LO potentials with chiral NN and 3N forces developed in Ref. [72]. These nonlocal potentials are called the sim potentials because they were simultaneously fit to scattering and bound-state observables in the pion-nucleon, NN, and few-nucleon sectors. Specifically, we use the two N2LO potentials with the NN and 3N momentum cutoffs $\Lambda=500$ (left panel) and $450$ MeV (right panel); both constrained by NN scattering data at laboratory energies $T_{\mathrm{lab}}\leqslant 290$ MeV [72]. To obtain the PNM energy per particle, we apply the Monte Carlo framework for MBPT nuclear matter calculations developed in Ref. [73]. The green bands in Fig. 2 depict the Hartree-Fock reference energy and the blue and red bands the MBPT results at second and third order, respectively. All MBPT calculations are performed at the normal-ordered two-body level. More details on MBPT for nuclear matter can be found in Ref. [48]. This MBPT framework uses the operatorial definition of the chiral potentials, allowing us to straightforwardly vary the values of physical constants, such as the pion mass, low-energy couplings, and more. Not all chiral potentials are given in their operatorial form; e.g., when evolved to lower resolution scales using renormalization group (RG) methods, their operatorial definition is usually unknown. For the solid lines in Fig. 2, we keep the values of the low-energy couplings at their best-fit values determined in Ref. [72] and vary only the pion mass, nucleon mass, and the pion decay constant. The $m_{\pi}$ dependence of the nucleon mass was discussed in Sec. II.1, and the $m_{\pi}$ dependence of the pion decay constant is given by
@@
with its value in the chiral limit $f=86.2$ MeV and the LEC $\bar{l}_{4}=4.3$ [58]. The pion-mass dependence of the axial coupling constant $g_{A}$ is relatively weak in the regime considered here, as can be seen in Figure 1 in Ref. [58], and thus we keep $g_{A}=1.29$ at the physical value. We use these potentials to identify overall trends and emphasize the need to explore the pion-mass dependence of a wider range of chiral potentials in the future.
@@
where $f(\theta)=\left(m_{\pi}(\theta)/m^{\rm phys}_{\pi}\right)^{2}$, $\eta=D_{2}(m^{\rm phys}_{\pi})^{2}/\tilde{C}_{{}^{1}S_{0}}^{\rm fit}$ and $\tilde{C}_{{}^{1}S_{0}}^{\rm fit}$ is the best-fit value in each isospin channel obtained in Ref. [72]. Here, $D_{2}$ is the LEC discussed in Ref. [66, 64] that captures the pion-mass dependence at NLO. We choose the two values $\eta=\pm 0.1$ to set the limits of this variation and adopt $m^{\rm phys}_{\pi}=138.04$ MeV for the physical pion mass. The dotted and dash-dotted lines in Fig. 2 depict the results for the PNM energy per particle with the modified LO coupling constant (19); the solid lines correspond to $\eta=0$ and thus $\tilde{C}_{{}^{1}S_{0}}(\theta)\equiv\tilde{C}_{{}^{1}S_{0}}$. Increasing the strength of $\tilde{C}_{{}^{1}S_{0}}(\theta)$, corresponding to $\eta=-0.1$, reduces $\Delta E(n_{\rm sat},m_{\pi})$, whereas decreasing the strength, corresponding to $\eta=+0.1$, increases $\Delta E(n_{\rm sat},m_{\pi})$.
Several important insights can be drawn from the results in Fig. 2. MBPT appears to converge more slowly with decreasing pion mass, which could be further investigated using a Weinberg eigenvalue analysis [74]. Although the leading Hartree-Fock energies indicate a systematic increase in interaction energy with $m_{\pi}$ across the range shown in Fig. 2, the second-order correction reverses this trend. The third-order correction is modest and of the same sign as the second-order correction. For relatively small reductions of the pion mass ($m_{\pi}\gtrsim 120$ MeV for $\Lambda=500$ MeV and $m_{\pi}\gtrsim 110$ MeV for $\Lambda=450$ MeV), the energy difference (17) is positive. The sensitivity of the results at the second and third order to the cutoff variation is larger than that of the LEC variation. These general trends are not altered by varying the momentum cutoff over the range shown in Fig. 2. However, the rather large dependence on the momentum cutoff is disconcerting and warrants a careful inclusion of the $m_{\pi}$ dependence of the short-range components of the force. As discussed earlier, we have studied the effect of the $m_{\pi}$ dependence of the leading LEC associated with scattering in the ${}^{1}S_{0}$ channel shows that the qualitative behavior is not changed. Based on these trends, we conclude that the interaction energy at $\theta=\pi$, where $m_{\pi}\approx 80$ MeV, can be smaller than that at $\theta=0$, where $m_{\pi}\approx 138$ MeV. However, the energy difference $\Delta E$ remains quite sensitive to the regularization scheme.
@@
III.3 Axion condensation in Relativistic Mean Field Theory
To derive concrete results for neutron stars, we use a model within RMFT (see Ref. [75] for a review). The Lagrangian of this theory is given by the following, omitting terms that will be discarded after taking the meson fields to have their mean field values.
@@ III.3 Axion condensation in Relativistic Mean Field Theory
$\mathcal{L}_{\sigma\omega\rho}$ contains all meson-meson couplings that are fit, along with the nucleon-meson couplings and meson masses, to known properties of nuclear matter. $\tau_{3}$ is the third Pauli matrix in isospin space. We use the IUFSU parameter set from Ref. [76]. The mean field values of the mesons and Fermi momenta of the nucleons and leptons are calculated by solving the Euler-Lagrange equations for each field and enforcing charge neutrality and $\beta$ equilibrium.
@@
For Scenario A, the simple case where only the nucleon masses are modified by finite $\theta$, Fig. 3 shows the energy per baryon, including axion energy and the reduction of the mass of the nucleons, of SNM, PNM, and beta equilibrated matter in this model. At fixed baryon density, the energy per baryon is minimized, leading to axion condensation due to the reduced nucleon mass. For small $\varepsilon$, the energy cost of having non-zero $\theta$ is small and axion condensation is favored at lower densities. Since the axions themselves contribute to the energy density, the energy per baryon of axion condensed matter will always diverge in the limit of zero density and a phase without axions will be favored. This has the effect of producing a minimum in the energy per baryon of axion condensed matter when $\varepsilon$ is sufficiently small and allows a stable zero pressure phase of nuclear matter at lower energy than normal nuclei. We will return to discuss objects made out of such matter in Sec. V.
@@
Of particular relevance to observations of neutron stars is the pressure at which neutron drip occurs. Neutron drip refers to the density at which it becomes energetically favorable for neutrons to come out of nuclei and form a background of free neutrons. As such, the neutron drip phase transition will occur when the baryon chemical potential is equal to the mass of the neutron (possibly modified by finite $\theta$) in the case where PNM is not self-bound. If the axion-condensed phase persists to the surface of the neutron star, but the negative axion pressure is larger than the pressure of electrons at neutron drip density (nuclei make a small contribution to the pressure), the neutron drip phase transition will not be found in the neutron star and the surface of the star will have a thin domain wall covering either a phase of nuclei in a dripped neutron background or bulk nuclear matter.
Not included thus far is the pion mass dependence of the nuclear force itself. The $\theta$-dependence of the masses of the $\sigma$ and $\rho$ mesons have been calculated in Refs. [56, 57]. They find that, comparing $\theta=0$ and $\theta=\pi$, the $\sigma$ mass decreases by about 6% and the $\rho$ mass decreases by about 2%. At low density, the RMFT EOS is dominated by the $\sigma$ meson and its modification is much larger than that of the vector mesons so we focus on the effect of modifying the scalar interaction. As a simple ansatz for modifications to the nuclear force, we modify the $\sigma$ mass with a tunable parameter $d_{\sigma}$ as follows:
@@
where $g_{\sigma}$ is a parameter of the RMFT model and $\chi$ is the volume fraction of nuclei when in a heterogeneous phase and unity in a homogeneous phase. At saturation density, $g_{\sigma}^{2}n_{\rm sat}/\sigma_{N}m_{\sigma}^{2}\simeq 11$ for the RMFT model we use so for Scenario B, axion condensation should be expected below $n_{\rm sat}$ for $\varepsilon\lesssim 0.8$ and in Scenario C, axion condensation should be expected below $n_{\rm sat}$, even for $\varepsilon=1$! The predictive power of this model is limited by the fact that we only consider isoscalar effects on the nuclear force. Axion condensation may also have an important impact on the symmetry energy of nuclear matter beyond the effect of $\Delta\sigma$. We expect this to have a more pronounced effect at high density where our model of the nucleon-axion interaction breaks down.
@@
Figure 4 shows the kinetic and interaction energy per baryon for Scenarios A, B, and C at $\theta=\pi$. Since this only includes the energy from the kinetic energy of baryons and nuclear forces, this quantity is independent of $\varepsilon$. Note that this differs from Fig. 3 in that the reduction of the nucleon masses is not included in this value.
When $\varepsilon$ is small enough that axion condensed matter is favored down to zero pressure, there can be striking effects on the crust of the star. We construct an axion-condensed inner crust of nuclei in a background of dripped neutrons, including Coulomb and surface energies. Since the constraints we derive on axion parameter space are tied to the properties of the crust, we choose a surface tension $\sigma=0.75\,\mbox{MeV}/\mbox{fm}^{2}$ tuned to recreate neutron drip pressure at $\theta=0$ calculated with more detailed microphysics (see, e.g., Ref. [77]). To be self-consistent between the two phases, the phase transition will be calculated between the RMFT model in all crust phases for all values of $\theta$ rather than using a tabulated crust EOS. Where relevant, we compare quantitative results to the calculation of crust properties calculation in the BPS and NV EOSs [77, 78].
@@
Figure 5 shows a phase diagram for the RMFT model as a function of pressure and $\varepsilon$. In the blue-shaded region, $\theta=\pi$ is favored while in the unshaded region, $\theta=0$ is favored. The cross-hatched regions contain the outer crust of nuclei and electrons and the diagonal-hatched regions contain the inner crust of nuclei, dripped neutrons, and electrons. Regions without hatching are composed of homogeneous nuclear matter. The diagonal colored lines show the domain wall pressure for an axion of a given mass with the specified $\varepsilon$ if the radius of the domain wall is 12 km. When the boundary region is treated as infinitely thin, this is the pressure difference across the domain wall. For small $\varepsilon$ where the axion condensed phase is favored for all densities, this acts as a cutoff pressure at the surface of the star; lower pressures can only exist in the finite size of the domain wall itself. Effects of the domain wall will be discussed in more detail in Sec. IV.2.
@@
Figure 5: Phase diagram for RMFT model with IUFSU parameter set and our three choices of $K_{\sigma}$. The colored diagonal lines show the domain wall pressure for a fiducial neutron star radius $R=12$ km and two example axion masses. The blue shading (no shading) indicates the pressures for which $\theta=\pi$ ($\theta=0$) is the favored phase. The hatching indicates the regions of increasing density within the neutron star: outer crust (cross-hatched), inner crust (diagonal hatched) and core (no hatch).
@@
As $K_{\sigma}$ is increased from 00, the pressure of the PNM surrounding the nuclei in the inner crust decreases significantly and, as a result, the chemical potential corresponding to the crust-core transition (the boundary between the diagonally hatched and unhatched regions) decreases. At the same time, a more attractive nuclear interaction causes nuclei to become more neutron-rich, increasing the electron chemical potential. An increased electron density corresponds to a larger pressure at neutron drip (the boundary between the diagonal hatched and cross-hatched regions). For sufficiently large $K_{\sigma}$, if $\varepsilon$ is small enough for an axion condensed crust to be present, there will be a phase transition directly from bulk nuclear matter to a phase with nuclei and free electrons, with no dripped neutrons. This occurs because bulk matter is favored when the chemical potential is less than $m_{n}(\theta=\pi)+\Gamma_{n}(\theta=\pi)$ where $\Gamma_{n}(\theta)$ is the binding of PNM in the case where neutron matter is self-bound (Scenarios B and C) and 00 in the case where it is not.
III.4 EOS of axion condensed matter